### v2

# Vidar: Embodied Video Diffusion Model for Generalist Bimanual Manipulation

Abstract
Bimanual robotic manipulation, which involves the coordinated control of two robotic arms, is foundational for solving challenging tasks. Despite recent progress in general-purpose manipulation, data scarcity and embodiment heterogeneity remain serious obstacles to further scaling up in bimanual settings. In this paper, we present Video Diffusion for Action Reasoning (Vidar), a two-stage framework that leverages large-scale, diffusion-based video pre-training and a novel masked inverse dynamics model for action prediction. We pre-train a general video diffusion model on 750K multi-view videos from three real-world bimanual robot platforms, utilizing a unified observation space that encodes robot, camera, task, and scene contexts. Our masked inverse dynamics model learns masks to extract action-relevant information from generated trajectories without requiring pixel-level labels, and the masks can effectively generalize to unseen backgrounds. Experimental results demonstrate that with only 20 minutes of human demonstrations on an unseen robot platform (only 1% of typical data requirements), Vidar generalizes to unseen tasks and backgrounds with strong semantic understanding, surpassing state-of-the-art methods. Our findings highlight the potential of video foundation models, coupled with masked action prediction, to enable scalable and generalizable robotic manipulation in diverse real-world settings.
Figure 1: Overview of Video Diffusion for Action Reasoning (Vidar), a video diffusion foundation model for video prediction and masked inverse dynamics model (MIDM) for action regression. Trained on 750K multi-view bimanual videos with test-time scaling (TTS) applied during testing, Vidar can adapt to new robot platforms with 20 minutes of demonstrations with state-of-the-art performance and generalize to unseen tasks with strong semantic understanding.
1 Introduction
Bimanual manipulation—the coordinated control of two robotic arms—is central to many high-precision tasks in both industrial automation and domestic robotics. In contrast to single-arm manipulation, dual-arm skills enable stable object handling, in-hand reorientation, and multi-point contact control, which are essential for tasks such as cloth folding, liquid pouring, or tool-assisted assembly. Recent work on vision-language-action (VLA) foundation models [1, 2, 3, 4] has demonstrated promising results by leveraging large-scale multimodal pre-training, but extending these models to general-purpose bimanual settings remains highly nontrivial.
The challenge lies in the inherent complexity of bimanual control: the action space scales combinatorially with the number of controllable joints, while task success demands tight temporal coordination between arms, precise contact dynamics, and long-horizon planning. These properties substantially increase the data requirements and sensitivity to platform-specific variations. In practice, progress is limited by two fundamental bottlenecks. First, data scarcity: collecting large-scale, high-quality bimanual demonstrations is labor-intensive, cost-expensive and hardware-specific, resulting in orders-of-magnitude less data than what is available for single-arm or vision-language domains. Second, embodiment heterogeneity: differences in robot morphology, control frequency and kinematics typically lead to poor cross-platform transferability, requiring extensive robot-specific fine-tuning to adapt pre-trained models [5, 6].
To address these limitations, we propose Video Diffusion for Action Reasoning (Vidar) to effectively leverage Internet-scale video data for generalist bimanual manipulation. As outlined in Figure 1, Vidar consists of two ky stages. It first re-trains a general video diffusion model with 750K multi-view bimanual videos collected from three real-world robot platforms, where embodiment heterogeneity is handled by a unified observation space. We further adopt test-time scaling [7] to enhance the performance of the foundation model when domain-specific data is scarce. The predicted videos are then decoded into actions by the Masked Inverse Dynamics Model (MIDM), which learns to focus on and robustly extract action-relevant regions in the generated frames, without requiring any pixel-level supervision.
Such two-stage design makes Vidar be able to fully explore the embodied data pyramid, which consists of three progressively structured levels: (1) Web Video Data, offering broad, unstructured task diversity from Internet-scale sources; (2) Heterogeneous Robotic Data, providing mid-level embodied videos across diverse robot morphologies; and (3) Embodied Human Demonstrations, delivering fine-grained, domain-specific examples. In particular, the web video data is used for vanilla pretraining of the video diffusion model, endowing the model with general semantic understanding and visual rendering capabilities; heterogeneous robotic data supports embodied pretraining, enhancing its grasp of robot-specific dynamics and manipulation tasks; and embodied human demonstrations are used for fine-tuning, allowing precise adaptation to the target domain. By harnessing the first two levels of the pyramid, Vidar achieves strong generalization with significantly greater domain data efficiency—often by orders of magnitude.
Empirically, Vidar achieves strong performance with only 20 minutes of human demonstrations (3 demonstrations per task on average) in a previously unseen robotic platform. Notably, the amount of embodied human demonstrations required by Vidar is remarkably low—only 1/81 of that used by RDT [3] (27 hours), and 1/1200 of that used by $\pi_{0.5}$ [8] (400 hours). Moreover, Vidar outperforms state-of-the-art baselines VPP [9] and UniPi [10] by 58% and 40% respectively. Also, it can generalize to novel tasks and backgrounds, with good semantic understanding.
Our main contributions are:
  • We present the first large-scale embodied video diffusion model for bimanual manipulation, trained on 750K diverse demonstrations across three robot platforms with a unified observation space.
  • We introduce a masked inverse dynamics model that learns to extract generalizable action-relevant features from generated videos without requiring segmentation labels.
  • We demonstrate effective few-shot generalization with only 20 minutes of target-domain demonstrations, achieving state-of-the-art performance across unseen tasks and backgrounds.
Our results suggest that combining video generation and masked control prediction offers a scalable path toward general-purpose robotic manipulation in the wild.
2 Method
Figure 2: We introduce a unified observation space that integrates multi-view robotic videos and language instructions. This enables the training of a video diffusion foundation model on 750K multi-view bimanual robotic episodes, allowing it to generalize to unseen robot platforms with only 20 minutes of human demonstrations. At inference time, it generates multiple candidate videos, then applies test-time scaling (TTS) to select the most suitable one. We also propose the masked inverse dynamics model (MIDM), which learns masks to identify and attend to action-relevant regions for robust frame-based action regression.
In this section, we formally define our problem and present Vidar in detail. The overview of our method is shown in Figure 2.
2.1 Problem Formulation
Let $\mathcal{L}$ denote the language instruction space, $\mathcal{O}$ the observation space, and $\mathcal{A}$ the action space. Our objective is to learn a conditional robot manipulation policy $\pi:\mathcal{L}\times\mathcal{O}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathbb{P}(\cdot)$ represents the set of all probability measures over the given space. Directly learning this policy poses significant challenges, mainly due to the large volume of demonstrations required to achieve generation and generalization over three different modalities: language, observation, and action. For example, RDT-1B [3] collects 27 hours of human demonstrations for adaptation to the Aloha robot, which greatly limits its applicability to new platforms.
Our key insight is to introduce a unified intermediate domain in which the gaps between robot platforms are reduced, while still preserving all the information necessary for action prediction. In this way, we form a two-stage method, consisting of intermediate state prediction and action prediction. In the first stage, pre-training over existing embodied datasets can be applied with greater transferability; in the second stage, action prediction is also easier since the input only contains one modality. Here we choose the video space $\mathcal{V}$ as the intermediate domain. In this way, we can decompose the policy: $\pi=G\circ I$, where $G:\mathcal{L}\times\mathcal{O}\rightarrow\mathbb{P}(\mathcal{V})$ is a video generation model and $I:\mathcal{V}\rightarrow\mathcal{A}$ is an inverse dynamics model.
Unified Observation Space.
By leveraging the video modality as an intermediate representation, data from heterogeneous embodiments can be mapped to a unified space. Denoting the number of cameras as $V$, we can define a unified observation space $\mathcal{U}\subseteq\mathcal{L}\times\mathcal{O}$:
$\mathcal{U}=\{\langle\mathbf{o},\mathbf{l}\rangle|\mathbf{o}=\mathrm{aggregate}(\mathbf{I}^{(1)},\mathbf{I}^{(2)},\cdots,\mathbf{I}^{(V)}),\mathbf{l}=l_{r}+l_{c}+l_{t},\},$ (1)
where $\mathrm{aggregate}(\mathbf{I}^{(1)},\mathbf{I}^{(2)},\cdots,\mathbf{I}^{(V)})$ are the aggregation of multi-view images (e.g., concatenation after proper down-sampling), and $l_{r}$, $l_{c}$, and $l_{t}$ are instructions related to the robotic platform, camera, and task, respectively. In this way, robot, camera, task, and multi-view information, instead of only task and single-view information, are provided as the condition for the video generation model, reducing the gap between pre-training and adaptation and enabling compositional generalization.
We choose the Aloha platform with three views as our bimanual robot in the target domain, since it is widely used and suitable for teleoperations [11, 3]. For our target robot, actions are defined as target joint/gripper positions without conditioning on the history, which means the inverse dynamics model can be viewed as sequential applications of a simpler mapping from one image to one action. With a little abuse of notation, we denote the simpler mapping as the inverse dynamics model throughout.
One important assumption underlying this approach is that the intermediate video modality contains all the information for action prediction. However, this assumption fails to hold for many robotic systems due to the specific platform setting, including the Aloha platform. In the pre-training part of Figure 2, we can see that the arm joints frequently fall outside the camera’s field of view, even with three different views available. In our target domain, we adjust the center camera to a new position where two arms can be fully captured, shown in the fine-tuning part of Figure 2. In this way, our robotic platform differs from any platform we encountered during pre-training, serving as an ideal testbed for showing adaptation capability with scarce data.
2.2 Video Generation Model
We adopt rectified flow [12, 13] models, which generate high-quality videos by modeling pixel flow over time. Specifically, the model parameterizes a flow function $v:\mathcal{V}\times\mathbb{R}\times\mathcal{L}\rightarrow\mathcal{V}$, which models the velocity of pixels as they transition from noisy video frames to target frames under certain conditions and time, leading to the following ODE over a video $x_{t}$ (with $x_{0}$ sampled from a Gaussian and $x_{1}$ as the output video):
$\frac{\mathop{}\!\mathrm{d}x_{t}}{\mathop{}\!\mathrm{d}t}=v(x_{t},t,c),\quad t\in[0,1].$ (2)
During training, we train the “velocity” $v$ to approximate the constant flow from $x_{0}$ to $x_{1}$. The flow matching loss is:
$L_{G}=\mathbb{E}_{c,t,x_{0},x_{1}}\left[\left\|(x_{1}-x_{0})-v(tx_{1}+(1-t)x_{0},t,c)\right\|^{2}\right].$ (3)
We also augment the model with an additional "unconditional" task $\emptyset$ and jointly train on randomly sampled video pairs to enable classifier-free guidance. During inference, we can use any condition $c$ in our unified observation space, and update the flow $v$ using classifier-free guidance:
$v^{\prime}(x_{t},t,c)=v(x_{t},t,c)+\omega(v(x_{t},t,c)-v(x_{t},t,\emptyset)),$ (4)
where $\omega$ is a hyperparameter controlling the guidance strength. This new flow $v^{\prime}$ is used to generate videos by solving the ODE described in Eq. 2.
Embodied Pre-training and Fine-tuning on Unified Observation Space.
To address the diversity and heterogeneity of real-world bimanual robots, we pre-train our video generation model on a large-scale corpus of approximately 750K open-sourced episodes, all projected into a unified observation space. This dataset includes episodes with three fixed camera views, a configuration common in bimanual setups [14, 15]. The primary challenge lies in harmonizing heterogeneous embodiments: different robots vary in camera placements, viewing angles, and sensor resolutions, leading to covariate shifts in perception. Moreover, wrist-mounted views often introduce task-irrelevant information, increasing input noise and model variance.
Unified Observation Construction. To mitigate this, we standardize each frame at timestep $t$ by resizing less informative views (e.g., wrist-mounted) to smaller dimensions, and concatenating them with primary views along the channel dimension. Formally, each observation $\mathbf{o}_{t}$ is constructed as $\mathbf{o}_{t}=\bigoplus_{k=1}^{V}\gamma_{k}\cdot\phi_{r_{k}}(\mathbf{I}_{t}^{(k)})$, where $\mathbf{I}_{t}^{(k)}$ is the RGB image from camera $k$, $\phi_{r_{k}}$ is a spatial resizing function, $\gamma_{k}\in[0,1]$ is a confidence weight for view $k$. This operation produces a consistent tensor shape across platforms and preserves both semantic and kinematic context. We also incorporate dataset-specific metadata indicating the robot and camera view type to support domain adaptation.
During pre-training, episodes are uniformly sampled and used to train a conditional video diffusion model to generate temporally coherent rollouts conditioned on initial observations. At fine-tuning time, we apply the same preprocessing pipeline and perform supervised fine-tuning (SFT) on all model parameters using a small number of human demonstration episodes collected from the target platform.
Test-time Scaling.
Diffusion-based video generation is inherently stochastic, leading to significant variance in the quality and task-relevance of sampled rollouts.
Naïvely sampling a single trajectory may result in incoherent or suboptimal generations, especially in cluttered or ambiguous scenes. To address this, we apply a test-time scaling strategy: given an observation prefix $\mathbf{o}_{1}$, we generate $K$ candidate video trajectories $\{\tilde{\mathbf{v}}^{(i)}_{1:T}\}_{i=1}^{K}$ using different random seeds. We then rank these trajectories using a pretrained evaluator (e.g., CLIP or a vision-language model) $q_{\eta}$ and select the highest-scoring one, i.e., $\arg\max_{i}\;q_{\eta}(\tilde{\mathbf{v}}^{(i)}_{1:T})$. This approach reduces sampling variance and consistently improves the quality of generated demonstrations with marginal computational overhead.
2.3 Masked Inverse Dynamics Model
Inverse dynamics models often suffer from poor generalization due to the presence of background noise, texture biases, and visual distractions in high-dimensional observations. Explicitly localizing action-relevant regions is difficult without dense annotations. We introduce a masked inverse dynamics model (MIDM) that learns to focus on task-relevant regions via implicit mask prediction. The model consists of two components: (1) a mask prediction network $U$ that outputs a spatial mask $m\in[0,1]^{H\times W}$ from an input frame $x$, and (2) an action regression network $R$ that predicts the action from the masked frame. Formally:
$m=U(x),\quad\hat{a}=R(\text{round}(m)\odot x),$
where $\odot$ denotes element-wise multiplication. The model is trained with the following loss:
$L_{I}=\mathbb{E}_{x,a}\left[l(\hat{a}-a)+\lambda\|m\|_{1}\right],$
where $l(\cdot)$ may be any suitable loss (e.g., MSE), and we train it using straight-through estimators. Here, the second term promotes spatial sparsity, encouraging the model to focus on minimal, task-critical regions without any segmentation supervision. This enables robust generalization to unseen environments and backgrounds, as shown in our experiments (Section 3).
3 Experiments
We want to prove the following hypotheses:
  1. H1:
    Vidar achieves superior success rates with only 20 minutes of target domain demonstrations;
  2. H2:
    Vidar generalizes effectively to unseen tasks and backgrounds;
  3. H3:
    Pre-training with a unified observation space benefits robot manipulation;
  4. H4:
    Masked inverse dynamics models exhibit greater generalization ability than the baseline.
3.1 Experimental Setup
3.1.1 Datasets
For pre-training, we utilize episodes sourced from Agibot-World [15], RoboMind [5], and RDT [3]. For Agibot-World, episodes are segmented into shorter clips using their frame-level annotations. As is required by the unified observation space, we provide descriptions of the robot type and camera placements in addition to the provided task instructions. The resulting pre-training dataset comprises 746,533 episodes.
For the target robot domain, we collect 20 minutes of human demonstration videos, covering 81 tasks across 232 episodes. This dataset is used to both fine-tune the video diffusion model and to train the masked inverse dynamics model. For more details of our datasets, see Appendix A.
3.1.2 Training and Inference
We utilize Vidu 2.0 [16] as our video generation model. We train for 23,000 iterations with a batch size of 128, where the first 10,000 steps involve pre-training on the diverse robotics dataset, followed by 13,000 fine-tuning steps. To reduce inference costs, we uniformly downsample the training videos to 8 frames per second (fps). For the masked inverse dynamics model, we adopt the U-Net [17] structure as the mask prediction network and the ResNet [18] structure as the action regression network, with $\lambda=3\times 10^{-3}$ (effects of $\lambda$ are shown in Appendix C). We train it exclusively on the fine-tuning dataset. We use AdamW [19] for all our training.
We use open-loop control for Vidar; the videos are generated in a single batch, without subsequent generation after the initial run. Using 8 NVIDIA Ampere-series 80GB GPUs, generating one video with 60 frames (7.5 seconds duration at 8fps) costs about 25 seconds. For test-time scaling, we choose $K=3$, and three videos with different random seeds are generated in parallel, evaluated by GPT-4o [20]. More details of training and inference can be found in Appendix B.
We acknowledge that Vidu 2.0 is not open source. To demonstrate the generality of our approach, we also perform small-scale experiments using the open-source HunyuanVideo model [21]. Related results are shown in Appendix E.
3.1.3 Baselines
We perform preliminary experiments over multiple baselines and find that adaptation with only 20 minutes of videos and about 3 demonstrations per task is too challenging for vision-language-action models. Thus, we choose two baselines that also incorporate video-level prior knowledge: UniPi and VPP. To ensure fair comparison, we reproduce these methods over the advanced Vidu 2.0 model.
UniPi.
We fine-tune the Vidu 2.0 model directly on our demonstration data, without any additional pre-training on our large robotics dataset. Additionally, we train an inverse dynamics model using a ResNet architecture.
VPP.
We use the same checkpoint for the video generation model as our method, and train a diffusion model to generate short action sequences based on the latent features during one-step video generation and CLIP [22] embeddings of the task instructions. Notably, they use closed-loop control, which means new action sequences are generated and executed after previous executions.
3.2 Experimental Results
H1: Success Rates.
We test two baseline methods and our method under three scenarios: seen task and background (six tasks), unseen task (five tasks), and unseen background (six tasks). Detailed explanations are as follows:
  • Seen Tasks & Backgrounds: three pick-and-place tasks (e.g., grasp tomato using left arm), two daily-life tasks (e.g., flip dice using right arm), and one bimanual task (lift the basket). The background we use is a cluttered office desk setup, featuring several computers situated behind the workspace.
  • Unseen Tasks: three daily-life tasks (e.g., stack the bowl on the steamer using the left arm) and two semantic tasks (e.g., grasp the shortest bread using the left arm).
  • Unseen Backgrounds: two pick-and-place tasks (e.g., grasp a tomato using the left arm) and four daily-life tasks (e.g., flip a dice using the left arm). For unseen backgrounds, we include a studio setup with a typical green screen and a daily workspace setting with two cupboards containing robot supplies, which exhibit reflective surfaces.
The success rates are shown in Table 1. We find that our method achieves superior success rates over all three scenarios, demonstrating the effectiveness of our method. More details can be found in Appendix C.
Table 1: Success rates of different methods and configurations over robot manipulation tasks. Vidar achieves high success rates across all three scenarios, with great generalization ability to unseen tasks and backgrounds.
Methods Seen Tasks & Backgrounds Unseen Tasks Unseen Backgrounds
VPP 4.5% 13.3% 0.0%
UniPi 36.4% 6.7% 22.2%
Vidar (Ours) 68.2% 66.7% 55.6%
H2: Generalization Ability.
Figure 3: Videos of predictions (left) and corresponding executions (right) of Vidar for challenging tasks. It can handle unseen tasks and unseen backgrounds with strong semantic understanding.
Quantitative results of our superior generation ability are shown in unseen scenarios of Table 1. We also provide some demonstrations in Figure 3, where Vidar demonstrates robust generalization to unseen tasks and backgrounds with strong semantic understanding. In Appendix D, we present more visualizations, including failure cases.
H3: Effectiveness of Pre-training.
We evaluate the video generation quality of the original Vidu 2.0 model and our pre-trained version over the unseen target domain using VBench [23]. As is shown in Table 2, we find that pre-training using large-scale robotic videos under our unified observation space enhances both the consistency and quality of generated frames, which are important for robot control tasks.
Table 2: VBench video quality measurements for different video model configurations in the unseen target domain. Embodied pre-training over the unified observation space benefits video generation.
Configurations Subject Consistency Background Consistency Imaging Quality
Vidu 2.0 0.565 0.800 0.345
+ Embodied Pre-training 0.855 0.909 0.667
H4: Effectiveness of MIDM.
Based on empirical observations of the Aloha robot’s error tolerances, we define a successful prediction as having a maximum infinity norm error of less than 0.06 for joint positions and less than 0.6 for gripper positions, for both arms. Using this criterion, we evaluate the success rates of our masked inverse dynamics model (MIDM) and a ResNet baseline on both training and testing sets. The results, presented in Table 3, show that our MIDM demonstrates superior generalization compared to the baseline. Additionally, examples of the learned masks are provided in Figure 4. Without any additional supervision, MIDM effectively captures action-relevant features and generalizes well to unseen backgrounds.
Table 3: Training and testing success rates and testing $1_{1}$ errors of different inverse dynamics models. Both the baseline and MIDM achieve high performances during training, but MIDM generalizes better during test time.
Inverse Dynamics Models Training Accuracy Testing Accuracy Testing $l_{1}$ Error
ResNet 99.9% 24.3% 0.0430
MIDM (Ours) 99.9% 49.0% 0.0308
Figure 4: Input images and corresponding masked images learned by the masked inverse dynamic model (MIDM). The two cases at the top are from a seen background, while the other two are from an unseen background. The predicted masks focus on the essential parts of robotic arms and have great generalization ability.
3.3 Ablation Study
We conduct an ablation study of our method by evaluating success rates on the same tasks presented in Table 1. The results are summarized in Table 4, where "w/o MIDM" means using the ResNet baseline. We find that both masked inverse dynamics models and test-time scaling are beneficial to the success rates.
Table 4: Ablation study of Vidar, where "w/o MIDM" means using the ResNet baseline. In three scenarios, both masked inverse dynamics models and test-time scaling are beneficial to the success rates.
Configurations Seen Tasks & Backgrounds Unseen Tasks Unseen Backgrounds
Vidar w/o TTS 45.5% 33.3% 44.4%
Vidar w/o MIDM 59.1% 26.7% 22.2%
Vidar 68.2% 66.7% 55.6%
4 Related Work
Vision-Language-Action Models (VLAs).
Vision-Language-Action (VLA) models integrate perception, language understanding, and action generation to enable general-purpose embodied intelligence. However, existing VLA systems rely heavily on large-scale, task-specific datasets—often requiring hundreds of thousands of trajectories—to achieve robust performance. For instance, RT-1 [24] uses 130K real-world episodes, while RT-2 [25] scales to 1B image-text pairs. Recent works like OpenVLA [2], Octo [26], Pi0 [27], and RDT-1B [3] further expand to millions of demonstrations across diverse embodiments. Despite these efforts, such data-intensive pipelines present a scalability bottleneck and limit generalization to novel tasks and domains, motivating more data-efficient and transferable alternatives.
Video Generation Models for Embodied AI.
Video generation models for embodied AI predict future scene dynamics to assist robotic planning and policy learning. Current works such as UniPi [28], RoboDreamer [29], Gen2Act [30], and GR-1 [31] mainly adopt text-conditioned video generation with a fixed single-camera view, demonstrating how a single-arm robot physically compliantly executes a series of actions. Additionally, Genie 2 [32] generates interactive 3D environments via autoregressive latent diffusion, enabling scalable training for embodied agents. These models implicitly encode physical laws through video synthesis, reducing reliance on real-world robotics data. However, current works mostly remain limited to single-arm robots and rely on a single camera view, which restricts their applicability in more complex real-world scenarios.
Coupled Video Generation and Action Synthesis.
Recent efforts couple video generation with action synthesis to enhance physical consistency and interoperability. For instance, Video-Prediction-Policy (VPP) [9] learns implicit inverse dynamics conditioned on future representations predicted by video diffusion model, while UVA [33] unifies video-action latent spaces with lightweight diffusion heads. Other works, such as VidMan [34] and Dreamitate [35], integrate low-level action representations with video prediction to predict actions. These approaches bridge spatiotemporal dynamics between vision and action, offering novel solutions for complex manipulation tasks. However, these approaches require end-to-end joint training of video generation and action prediction models, which limits their flexibility and adaptability. Integrating a more advanced video generation model would necessitate retraining the entire system.
5 Conclusion and Broader Impact
We presented Vidar, a generalizable framework for bimanual robotic manipulation that addresses the core challenges of data scarcity and embodiment heterogeneity. By combining large-scale, diffusion-based video pretraining over a unified observation space with a masked inverse dynamics model, Vidar enables accurate action prediction from multi-view visual observations and language instructions, requiring only minimal demonstrations in new environments. Our experiments demonstrate that Vidar consistently outperforms existing methods and exhibits strong generalization to unseen tasks and backgrounds, highlighting its capacity for semantic understanding and transfer.
Appendix A Dataset Details
The details of our datasets are presented in Table 5, with corresponding visualizations shown in Figure 5. We filter out episodes with fewer than three views or shorter than four seconds, and we further segment lengthy episodes in the Agibot dataset. For each dataset, we provide information about the associated robot and camera setups, which form part of the unified observation space. During pre-training, sampling ratios for each dataset are set proportionally to their sizes. It is worth noting that our Vidar dataset differs from all pre-training datasets in these aspects, and we additionally leverage GPT-4o to augment its task annotations.
Table 5: Detailed information about datasets. Dataset instructions correspond to the robot and camera components of the unified observation space (see Figure 2).
Dataset Size Dataset Instructions
Agibot 726,843 The whole scene is in a realistic, industrial art style with three views: a fixed high camera, a movable left arm camera, and a movable right arm camera. The genie-1 robot is currently performing the following task:
RDT 6,023 The whole scene is in a realistic, industrial art style with three views: a fixed front camera, a movable left arm camera, and a movable right arm camera. The aloha robot is currently performing the following task:
RoboMind Franka 6,395 The whole scene is in a realistic, industrial art style with three views: a fixed camera on the opposite side, a fixed left camera, and a fixed right camera. The franka robot is currently performing the following task:
RoboMind Aloha 7,272 The whole scene is in a realistic, industrial art style with three views: a fixed front camera, a movable left arm camera, and a movable right arm camera. The aloha robot is currently performing the following task:
Vidar 232 The whole scene is in a realistic, industrial art style with three views: a fixed rear camera, a movable left arm camera, and a movable right arm camera. The aloha robot is currently performing the following task:
Figure 5: Visualizations of datasets. The pre-training datasets feature a variety of robots and camera configurations, whereas our Vidar dataset distinctively differs from all of these settings.
Appendix B Training and Inference Details
Using 64 NVIDIA Ampere-series 80GB GPUs, we train Vidu 2.0 for 23,000 iterations (10,000 for pre-training and the remainder for fine-tuning), which takes about 64 hours. Note that for all fine-tuning procedures, we employ full-parameter fine-tuning. For MIDM, we use 8 NVIDIA Hopper-series 80GB GPUs for 60,000 training iterations, taking about 5 hours. Additional MIDM parameters are provided in Table 6.
During inference, the video diffusion models are deployed in the cloud, while only the lightweight MIDM is executed locally. For test-time scaling, we uniformly sample 5 – 7 frames from the generated videos and use GPT-4o to select the best result.
Table 6: Parameters of MIDM.
Parameters Values
Number of Parameters 92 Million
U-Net Down-sampling/Up-sampling Layers 5
ResNet Structure ResNet-50
Action Prediction Loss Smoothed $l_{1}$ Loss
Learning Rate $2\times 10^{-5}$
Warm-up 6000 Steps
AdamW $\beta$ (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 )
AdamW $\epsilon$ $10^{-8}$
AdamW Weight Decay $10^{-2}$
Appendix C Additional Experimental Results
A more detailed version of Table 1 is provided in Table 7. We also investigate the impact of the parameter $\lambda$ in MIDM, with results summarized in Table 8 and illustrated in Figure 6. Notably, $\lambda=3\times 10^{-3}$ yields the best performance.
Table 7: Detailed success rates of various methods and configurations on robot manipulation tasks. "L," "R," and "B" indicate the use of the left arm, right arm, or both arms, respectively. "w/o TTS" and "w/o MIDM" correspond to the ablation settings described in Table 4.
Scenarios/Tasks Success Rates of Different Methods/Configurations
Seen Tasks & Backgrounds UniPi VPP Vidar w/o TTS w/o MIDM
Grasp the Tomato (L) 60.0% 0.0% 60.0% 60.0% 80.0%
Grasp the Tomato (R) 20.0% 0.0% 80.0% 60.0% 60.0%
Lift the Basket (B) 66.7% 0.0% 66.7% 33.3% 33.3%
Flip the Dice to Point One on the Top (L) 0.0% 33.3% 33.3% 0.0% 33.3%
Grab the Bottle (L) 0.0% 0.0% 66.7% 33.3% 33.3%
Get the Toast from the Toaster (L) 66.7% 0.0% 100.0% 66.7% 100.0%
Total 36.4% 4.5% 68.2% 45.5% 59.1%
Unseen Tasks UniPi VPP Vidar w/o TTS w/o MIDM
Place the Bowl on the Steamer (L) 0.0% 0.0% 66.7% 66.7% 33.3%
Grasp the Shortest Bread (L) 0.0% 0.0% 66.7% 33.3% 0.0%
Grasp the Shortest Bread (R) 0.0% 0.0% 66.7% 33.3% 33.3%
Wipe the Table with Rag (L) 33.3% 33.3% 66.7% 33.3% 66.7%
Wipe the Table with Rag (R) 0.0% 33.3% 66.7% 0.0% 0.0%
Total 6.7% 13.3% 66.7% 33.3% 26.7%
Unseen Backgrounds UniPi VPP Vidar w/o TTS w/o MIDM
Grasp the Tomato (L) 0.0% 0.0% 66.7% 33.3% 0.0%
Flip the Die to Point One on the Top (L) 0.0% 0.0% 33.3% 33.3% 0.0%
Flip the Die to Point One on the Top (R) 0.0% 0.0% 33.3% 0.0% 0.0%
Pick the Carrot on the Green Plate (L) 33.3% 0.0% 33.3% 66.7% 0.0%
Place the Bowl on the Plate (L) 33.3% 0.0% 66.7% 66.7% 33.3%
Place the Bowl on the Plate (R) 66.7% 0.0% 100.0% 66.7% 100.0%
Total 22.2% 0.0% 55.6% 44.4% 22.2%
(a) $\lambda=10^{-1}$
(b) $\lambda=10^{-2}$
(c) $\lambda=3\times 10^{-3}$
(d) $\lambda=10^{-3}$
(e) $\lambda=10^{-4}$
Figure 6: Masked images learned by the masked inverse dynamic model (MIDM) with different values of $\lambda$.
Table 8: Analysis of hyperparameter $\lambda$ in the masked inverse dynamics model. The success rates are robust over a large range, and $\lambda=3\times 10^{-3}$ achieves the best performance.
$\lambda$ Training Accuracy Testing Accuracy Testing $l_{1}$ Error
$10^{-1}$ 99.1% 7.1% 0.0670
$10^{-2}$ 99.8% 39.9% 0.0331
$3\times 10^{-3}$ 99.9% 49.0% 0.0308
$10^{-3}$ 99.9% 40.7% 0.0338
$10^{-4}$ 99.8% 24.4% 0.0461
Appendix D Additional Visualizations of Vidar
More demonstrations, including two failed cases, are shown in Figure 7.
Figure 7: Videos of predictions (left) and corresponding executions (right) of Vidar for more challenging tasks. Both successful and failed cases are presented.
Appendix E Reproduction using the HunyuanVideo Model
We also reproduced our results using the open-source HunyuanVideo Model. The training parameters are detailed in Table 9. Utilizing 64 NVIDIA Hopper-series 80GB GPUs, we trained the model for 12,000 iterations (10,000 for pre-training and the remainder for fine-tuning), which required approximately 54 hours.
Due to limited resources, we evaluated the model on six tasks. The results are presented in Table 10 and Figure 8. We plan to open-source our implementations related to both the HunyuanVideo Model and our MIDM.
Table 9: Parameters of the HunyuanVideo Model.
Parameters Values
Number of Parameters 13 Billion
Learning Rate (Pre-train) $1\times 10^{-4}$
Learning Rate (Fine-tune) $5\times 10^{-5}$
Warm-up 500 Steps
Optimizer AdamW
AdamW $\beta$ (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 )
AdamW $\epsilon$ $10^{-8}$
AdamW Weight Decay 0
Table 10: Success rates of Vidar reproduced using the HunyuanVideo Model.
Tasks Success Rates
Grasp the Apple (Seen Task) 75.0%
Grab the Bottle (Seen Task) 100.0%
Grasp the Cup by its Handle (Seen Task) 25.0%
Lift the Steamer (Unseen Task) 50.0%
Grasp the Apple and Put it into the Steamer (Unseen Task) 100.0%
Stack One Cube on Top of the Other Cube (Unseen Task) 20.0%
Total 58.3%
Figure 8: Videos of predictions (left) and corresponding executions (right) of Vidar reproduced using the HunyuanVideo model for challenging tasks.
Appendix F Hardware Details
Hardware details are shown in Figure 9 and Table 9.
Figure 9: Our robotic platform.
Table 11: Hardware Information.
Parameters Values
Degree of Freedom 2×(6+1)=142$\times$(6+1)=142 × ( 6 + 1 ) = 14
Cameras 3 RGB Cameras
[…]